OpenStudy (natsudragneel):

Create a problem that uses the associative property to find the sum of at least 3 rational numbers. Show how the associative property is used to solve the problem. WILL MEDAL AND FAN!!!!!!!!!

OpenStudy (natsudragneel):

Help please!! Make two different problems!!

OpenStudy (natsudragneel):

Will medal and fan!

OpenStudy (518nad):

"Definition: The associative property states that you can add or multiply regardless of how the numbers are grouped. By 'grouped' we mean 'how you use parenthesis'. In other words, if you are adding or multiplying it does not matter where you put the parenthesis. Add some parenthesis any where you like!." gogle

OpenStudy (natsudragneel):

I need two problems please! Just make up two problems that have 3 rational numbers using the associative property?

OpenStudy (callielovesyhuu):

2+8=8+2 4+3=3+4 1+8=8+1

OpenStudy (callielovesyhuu):

2+1+3=3+1+2 5+6+7=7+6+5

OpenStudy (natsudragneel):

Thank you!